# A new source of purely finite matricial fields

### 25 Mar 2026 v1

@@ Abstract.
Our work has several applications. Firstly, as a consequence of MF, the Brown–Douglas–Fillmore semigroups of many new reduced $C^{*}$-algebras are not groups. Secondly, we obtain that arbitrary graph products of residually finite exact MF (resp., PMF, PFF) groups are MF (resp., PMF, PFF), yielding a significant generalization of the breakthrough work of Magee–Thomas. Thirdly, our work resolves the open problem of proving PFF for 3-manifold groups. More generally all groups that virtually embed into RAAGs are PFF. Prior to our work, PFF was not known even in the case of free products. Our results are of geometric significance since PFF is the property that is used in Antoine Song’s approach in the theory of minimal surfaces.
@@ 1.1. Matricial field
In other words, $G$ is MF if there exists an ultrafilter $\mathcal{U}$ on $\mathbb{N}$ and a trace-preserving *-homomorphism from the reduced $C^{*}$-algebra, $\pi:C^{*}_{r}(G)\to\prod_{\mathcal{U}}\mathbb{M}_{n}(\mathbb{C})$, where the target $C^{*}$-algebra is the ultraproduct of matrix algebras with respect to the operator norm. Furthermore, $G$ is said to be purely MF (PMF) if the maps $u$ from above can be chosen as actual homomorphisms. $G$ is said to be a purely finite field (PFF) if the image of the homomorphisms are finite at each stage. $G$ is further said to be a purely permutation field (PPF) if in fact these homomorphisms are homomorphisms to finite permutation groups composed with standard irreducible representations of permutation groups. Note that the notations PMF and PPF are borrowed from [28]. The relationship between these notions is transparent and is given below:
@@
Outside of intrinsic interest in random matrix theory, these properties have phenomenal applications. MF has applications to operator algebras. If a $C^{*}$-algebra is MF but not quasidiagonal, then $A$ has an extension by the compact operators $\mathbb{K}$ which is not invertible in the sense of Brown, Douglas and Fillmore [6] (see for instance the remark at the end of Section 2 in [34]). In particular if a group is non amenable (hence not quasidiagonal [18]) and MF, then $Ext(C^{*}_{r}(G))$ is not a group. PMF has certain geometric applications [27], including spectral gaps of certain Laplacians acting on vector bundles. PFF has remarkably found use in the theory of minimal surfaces through work of A. Song in his approach [38]. PFF is also interestingly in the spirit of soficity [30] for algebras: it is not just a finite-dimensional approximation of the reduced group $C^{*}$-algebra, but even a finite approximation. PPF is the most desirable property here not only because it encompasses all of the above, but also because it plays a fundamental role in deep applications to the study of optimal spectral gaps of graphs and hyperbolic manifolds [5, 22]. PPF is however quite difficult to access, especially because it fails in certain simple examples (see the discussion before Remark 1.5). The main novelty of the present article is that despite this it is still possible, in significant generality, to bypass PPF and access PFF instead. For further details and references concerning these notions we point the reader to the beautiful surveys [28, 40].
@@
Importantly, these properties are much harder to prove than weaker representability properties such as soficity or Connes embeddability [30], because it involves establishing strong convergence. Despite the vast collection of examples of soficity and hyperlinearity, at the moment only a handful of examples of groups are known to be MF/PMF/PFF/PPF. Moreover, it is noteworthy to point out that proving such properties in each of these cases has been profoundly difficult and needed deep insights. The current list of examples includes MF for amenable groups [39]; PPF for free groups [5]; PPF for limit groups [26]; PMF for right angled Artin groups (RAAGs) [27]; MF for crossed products by free groups acting on amenable groups [31]; MF for $G_{1}*_{H}G_{2}$ where $G_{1},G_{2}$ are amenable and $H$ is a normal subgroup [33]. These properties are easily seen to be closed under taking subgroups and MF/PMF/PFF are closed under taking finite index overgroups (Lemma 7.1 in [26], see also Lemma 2.4). Being closed under finite index overgroups seems open in the case of PPF. It is currently known that MF/PMF is also closed under free products [10, 20]. Moreover, MF/PMF/PFF is also closed under direct products provided one of the groups is exact [17]. Strikingly, PPF is necessarily not closed under direct products, see [28] for a proof that $\mathbb{F}_{2}\times\mathbb{F}_{2}\times\mathbb{F}_{2}$ is not PPF. Despite all of these results, the situation for wider natural families of groups, in particular for amalgamated free products, has remained a challenging open problem.
@@ 1.2. Main results
In this paper we are able to make significant new progress and expand on the collection MF and notably also PMF and PFF groups by developing a new $C^{*}$-algebraic approach to the problem. Our main result is the following:
@@
Suppose $G$ is an MF (resp., PMF, PFF) group and $H<G$ is separable (i.e., $H=\cap_{i\in\mathbb{N}}H_{i}$ where $H_{i}<G$ are finite index subgroups). Let $L$ be a residually finite MF (resp., PMF, PFF) group such that either $G$ or $L$ is exact. Then the amalgamated free product $G*_{H}(H\times L)$ is MF (resp., PMF, PFF).
Our approach is inspired by recent progress on selflessness which is a form of intrinsic $C^{*}$-free independence in ultrapowers [32, 2, 29], in particular the work of Ozawa (see Section 4.6 of [7]). In order to demonstrate the power of our result, we assemble a few corollaries. Firstly, we observe that there is an isomorphism $G\ast_{H}(H\times\mathbb{Z})\cong\ast_{H}G\rtimes\mathbb{Z}$ where $_{H}G$ denotes the infinite group double of $G$ over $H$, indexed by elements of $\mathbb{Z}$, and $\mathbb{Z}$ acts on it by permuting the copies of $G$ in accordance with the left multiplication action of $\mathbb{Z}$ on itself. We also recall that, if $G$ and $H$ are PFF and one of them is exact, then $G\times H$ is PFF (Lemma 2.4). PFF groups are also automatically residually finite. This yields:
@@
We point out that strikingly even the case of free products remained open for the PFF property until our work. Outside of families of LERF groups (see a list of such examples in [15]), a nice instance of such inclusions $H<G$ which can be fed into our Corollary 1.2 is quasi-convex subgroups of cubulated hyperbolic groups. Indeed, cubulated hyperbolic groups are subgroups of RAAGs [1] and hence PFF (our Corollary 1.3 below), and quasi-convex subgroups are separable [19]. We also point out that even soficity or Connes embeddability is unknown beyond the generality we consider in such amalgamated free products.
Another natural setting for separability is in the world of graph products [23].
@@
Our result above offers a significant generalization of the breakthrough work of Magee–Thomas [27]. We emphasize that even in the case of RAAGs, our proof is different. While we still use the MF/PMF/PFF property of $\mathbb{F}_{n}$ [16, 5] as a black box in our proof, our approach avoids the use of other random matrix theory results [9], as well as the Crisp–Laca universality theorem [11] which are both crucial in [27]. More importantly, the reach of our results is higher and can handle the PFF property for RAAGs, which has been an open problem in the field [^1] We thank Ramon van Handel for sharing with us the statement of this open problem.. This particular result is of significance because this is exactly the ingredient that goes into the breakthrough work of A. Song on minimal surfaces in Euclidean unit spheres [38]. A. Song’s paper accesses PFF via the PPF result of Bordenave–Collins [5], and the follow up paper [3] accesses PFF for surface groups via Louder–Magee’s PPF for surface groups [26]. However, our results open the door to vast new families of groups, without having to go through PPF. This in particular includes 3-manifold groups and a plethora of other examples that virtually embed into RAAGs [1, 19]. We document this below.
@@
Fundamental groups of closed 3-manifolds are PFF. Then $G$ is PFF. More generally, all groups that virtually embed into RAAGs are PFF.
@@
We now discuss optimality aspects of our results. First note that the exactness is a rather essential ingredient, in for instance Corollary 1.3. Indeed even in the tensor product case it is not known if MF is preserved in full generality [17]. A natural question arises of whether our results can extend to accessing PPF. We point out that this is in fact not possible. For instance, our Corollary 1.3 simply cannot generalize to PPF as stated because as mentioned earlier, Magee has shown that the group $\mathbb{F}_{2}\times\mathbb{F}_{2}\times\mathbb{F}_{2}$ is not PPF [28], and this is indeed a RAAG. For this reason PFF seems to be the best one can hope for in these general families of groups. The challenge with PPF is that it is not amenable to being stable under direct products (or finite index overgroups), due to the following fundamental issue: tensoring does not preserve the image being in the orthogonal complement of the invariant vectors for permutation matrices.
Remark 1.5.
@@
We conclude with a minor remark that our work also provides an alternative proof of PFF for limit groups also (see Remark 2.6). It is important to clarify however that the work [26] proves PPF for limit groups and this cannot be addressed currently by our method. The proof of this in our work combines our Theorem 1.1 and the fact that limit groups arise as iterated extensions of centralizers (Theorem 4.2 of [8]), and Hall’s theorem for limit groups [45]. We also point out that the proof of Corollary 1.3 follows from the fact that every graph product is constructed in an iterated fashion using the amalgamated free products in Theorem 1.1, and additionally the amalgams are separable because they are retracts (Lemma 3.9 of [23]). We also remark that our approach might also naturally be applied to get more examples of MF/PMF/PFF groups among the family of graph wreath products [14].
@@ 1.3. Acknowledgments
Open Access and Data Statement
@@ 2. Proof of main result
Let $B\subset A$ be an inclusion of $C^{*}$-algebras, and $E:A\to B$ be a faithful conditional expectation. Then define the Toeplitz algebra of $E$ to be the universal $C^{*}$-algebra
@@
$T$ is called the creation operator.
The following Theorem 2.2 is due to Shlyakhtenko in [36] in the von Neumann algebra setting (see also [13]). The proof in the $C^{*}$-setting is more or less identical. Nevertheless, we provide a proof for the convenience of the reader (following [7] wherein this result also appears). Our inspiration for employing this result in our context comes from the recent outburst of developments surrounding selfless $C^{*}$-algebras [32, 2], especially the proof technique of Ozawa in [29] and also the upcoming key work [25]. The second and fourth authors are grateful to M. Junge and L. Robert for the collaboration on [25].
Before stating and proving Theorem 2.2, we emphasize that the beautiful gauge-invariant uniqueness theorem (this is explicitly [7, 4.6.18]), is used in the proof of Theorem 2.2. We highly encourage the reader to look at Section 4.6 in the treatise [7] which offers a clean treatment of these aspects.
@@
First consider $A\ast_{B}(B\otimes\mathcal{T})$ where $\mathcal{T}$ is the Toeplitz algebra. Let $S$ be the generating isometry in $\mathcal{T}$. Then for $a\in A$, $S^{*}aS=E(a)$. Hence there is a map $\pi:\mathcal{T}(E)\to A\ast_{B}(B\otimes\mathcal{T})$ which restricts to identity on $A$ and sends $T\mapsto S$. This map is surjective as $A\ast_{B}(B\otimes\mathcal{T})$ is generated by $A$ and $\mathcal{T}$.
We will prove $\pi$ is injective using the gauge-invariant uniqueness theorem for Toeplitz-Pimsner algebras [7, 4.6.18]. For this, note that the canonical conditional expectation $E^{\prime}:A\ast_{B}(B\otimes\mathcal{T})\to A$ acts by $\operatorname{id}\otimes\omega$ on the second factor, where $\omega$ is the vacuum state on the Toeplitz algebra. So we have
@@
where we used that $S$ commutes with $E(a)\in B$. So $\pi(A)\cap\overline{\operatorname{span}}\{a_{1}Sb_{1}b_{2}^{*}S^{*}a_{2}^{*}:a_{1},b_{1},b_{2},a_{2}\in A\}=\{0\}$, and so $\pi$ satisfies the hypothesis of the gauge-invariant uniqueness theorem.
@@
$G\ast_{H}(H\times\mathbb{Z})\to[G/K_{i}\ast_{H_{i}/K_{i}}(H_{i}/K_{i}\times\mathbb{Z})]\times G.$
@@
Let $E:C_{r}^{*}(G)\to C_{r}^{*}(H)$ and $E_{i}:C_{r}^{*}(G/K_{i})\to C_{r}^{*}(H_{i}/K_{i})$ be the respective conditional expectations. Let $q_{i}:\mathbb{C}[G]\to C_{r}^{*}(G/K_{i})$ be the map induced from the quotient map $G\to G/K_{i}$. Denote by $T$ the creation operator in $\mathcal{T}(E)$ and by $T_{i}$ the creation operator in $\mathcal{T}(E_{i})$.
Take $\mathbb{C}[G]\to\prod_{\mathcal{U}}\mathcal{T}(E_{i})\otimes C_{r}^{*}(G)$ to be the map that sends $\lambda_{g}$ to $(q_{i}(\lambda_{g})\otimes\lambda_{g})_{\mathcal{U}}$ for $g\in G$, where we recall $\mathcal{T}(E_{i})$ contains a copy of $C_{r}^{*}(G/K_{i})$. By Fell’s absorption principle, this is isometric if $\mathbb{C}[G]$ is given the reduced norm, and so extends to an embedding $C_{r}^{*}(G)\hookrightarrow\prod_{\mathcal{U}}\mathcal{T}(E_{i})\otimes C_{r}^{*}(G)$. Identifying the image of this embedding with $C_{r}^{*}(G)$, it is easy to check that the conditional expectation $E$ under this identification sends $(q_{i}(\lambda_{g})\otimes\lambda_{g})_{\mathcal{U}}$ to $(q_{i}(E(\lambda_{g}))\otimes\lambda_{g})_{\mathcal{U}}$ for $g\in G$. Now note
$(T_{i}^{*}q_{i}(\lambda_{g})T_{i}\otimes\lambda_{g})_{\mathcal{U}}=(E_{i}(q_{i}(\lambda_{g}))\otimes\lambda_{g})_{\mathcal{U}}=(q_{i}(E(\lambda_{g}))\otimes\lambda_{g})_{\mathcal{U}}.$
The latter equality can be verified easily using the fact that $H_{i}$ decreases to $H$. Hence, the isometry $(T_{i}\otimes 1)_{\mathcal{U}}$ satisfies the universal property of the creation operator in $\mathcal{T}(E)$ and this gives a *-homomorphism
$\mathcal{T}(E)\to\prod_{\mathcal{U}}\mathcal{T}(E_{i})\otimes C_{r}^{*}(G).$
This map is in fact an embedding. This follows from the gauge-invariant uniqueness theorem for Toeplitz–Pimsner algebras [7, Theorem 4.6.18]. Indeed, note that $(E_{i}^{\prime}\otimes\operatorname{id}_{C_{r}^{*}(G)})_{\mathcal{U}}$ defines a conditional expectation from $\prod_{\mathcal{U}}\mathcal{T}(E_{i})\otimes C_{r}^{*}(G)$ onto a subalgebra containing the image of $C_{r}^{*}(G)$, where $E_{i}^{\prime}:\mathcal{T}(E_{i})\to C_{r}^{*}(G/K_{i})$ are the canonical conditional expectations. The condition needed for the gauge-invariant uniqueness theorem easily follows, as in the proof of Theorem 2.2.
Now, by Theorem 2.2, restricting to $C^{*}(G,\frac{1}{2}(T+T^{*}))\subset\mathcal{T}(E)$ yields an embedding
$\begin{split}&C_{r}^{*}(G)\ast_{C_{r}^{*}(H)}(C_{r}^{*}(H)\otimes C[-1,1])\\ \hookrightarrow&\prod_{\mathcal{U}}[C_{r}^{*}(G/K_{i})\ast_{C_{r}^{*}(H_{i}/K_{i})}(C_{r}^{*}(H_{i}/K_{i})\otimes C[-1,1])]\otimes C_{r}^{*}(G).\end{split}$
That this is trace-preserving can be easily verified. Since $\frac{1}{2}(T+T^{\ast})$ is sent to $(\frac{1}{2}(T_{i}+T_{i}^{\ast}))_{\mathcal{U}}$, the map acts as the identity map on C[1,1]C[-1,1], so the result now follows by noting $C_{r}^{*}(\mathbb{Z})$ embeds into C[1,1]C[-1,1] trace-preservingly.
For the second part, we note that the map is indeed induced by the sequence of homomorphisms

### 30 Mar 2026 v3

@@ Abstract.
Our work has several applications. Firstly, as a consequence of MF, the Brown–Douglas–Fillmore semigroups of many new reduced $C^{*}$-algebras are not groups. Secondly, we obtain that arbitrary graph products of residually finite exact MF (resp., PMF, PFF) groups are MF (resp., PMF, PFF), yielding a significant generalization of the breakthrough work of Magee–Thomas. Thirdly, our work resolves the open problem of proving PFF for fundamental groups of closed hyperbolic 3-manifolds. More generally all groups that virtually embed into RAAGs are PFF. Prior to our work, PFF was not known even in the case of free products. Our results are of geometric significance since PFF is the property that is used in Antoine Song’s approach towards the existence of minimal surfaces of negative curvature in round spheres.
@@ 1.1. Matricial field
In other words, $G$ is MF if there exists an ultrafilter $\mathcal{U}$ on $\mathbb{N}$ and a trace-preserving *-homomorphism from the reduced $C^{*}$-algebra, $\pi:C^{*}_{r}(G)\to\prod_{\mathcal{U}}\mathbb{M}_{n}(\mathbb{C})$, where the target $C^{*}$-algebra is the ultraproduct of matrix algebras with respect to the operator norm. Furthermore, $G$ is said to be purely MF (PMF) if the maps $u$ from above can be chosen as actual homomorphisms. $G$ is said to be a purely finite field (PFF) if additionally the image of the homomorphisms are finite at each stage. $G$ is further said to be a purely permutation field (PPF) if in fact these homomorphisms are homomorphisms to finite permutation groups composed with standard irreducible representations of permutation groups. Note that the notations PMF and PPF are borrowed from [37]. The relationship between these notions is transparent and is given below:
@@
Outside of intrinsic interest in random matrix theory, these properties have phenomenal applications. MF has applications to operator algebras. If a $C^{*}$-algebra is MF but not quasidiagonal, then $A$ has an extension by the compact operators $\mathbb{K}$ which is not invertible in the sense of Brown, Douglas and Fillmore [9] (see for instance the remark at the end of Section 2 in [44]). In particular if a group is non amenable (hence not quasidiagonal [22]) and MF, then $Ext(C^{*}_{r}(G))$ is not a group. PMF has certain geometric applications [36], including spectral gaps of certain Laplacians acting on vector bundles. PFF has remarkably found use in the theory of minimal surfaces through work of A. Song in his approach [48]. PFF is also interestingly in the spirit of soficity [39] for algebras: it is not just a finite-dimensional approximation of the reduced group $C^{*}$-algebra, but even a finite approximation. PPF is the most desirable property here not only because it encompasses all of the above, but also because it plays a fundamental role in deep applications to the study of optimal spectral gaps of graphs and hyperbolic manifolds [7, 28]. PPF is however quite difficult to access, especially because it fails in certain simple examples (see Subsection 1). One of the main novelties of the present article is that, despite this, we demonstrate that it is still possible in significant generality to bypass PPF and access PFF instead. For further details and references concerning these notions we point the reader to the beautiful surveys [37, 50].
@@
Importantly, these properties are much harder to prove than weaker representability properties such as soficity or Connes embeddability [39], because it involves establishing strong convergence. Despite the vast collection of examples of soficity and hyperlinearity, at the moment only a handful of examples of groups are known to be MF/PMF/PFF/PPF. Moreover, it is noteworthy to point out that proving such properties in each of these cases has been profoundly difficult and needed deep insights. The current list of examples includes MF for amenable groups [49]; PPF for free groups [7]; PPF for limit groups [34]; PMF for right angled Artin groups (RAAGs) [36]; MF for crossed products by free groups acting on amenable groups [40]; MF for $G_{1}*_{H}G_{2}$ where $G_{1},G_{2}$ are amenable and $H$ is a normal subgroup [43]. These properties are easily seen to be closed under taking subgroups and MF/PMF/PFF are closed under taking finite index overgroups (Lemma 7.1 in [34], see also Lemma 2.4). Being closed under finite index overgroups seems open in the case of PPF. It is currently known that MF/PMF is also closed under free products [12, 25]. Moreover, MF/PMF/PFF is also closed under direct products provided one of the groups is exact [21]. Strikingly, PPF is necessarily not closed under direct products, see [37] for a proof that $\mathbb{F}_{2}\times\mathbb{F}_{2}\times\mathbb{F}_{2}$ is not PPF. Interestingly, it is also known that $SL_{4}(\mathbb{Z})$ is not PMF [35]. Despite all of these results, the situation for wider natural families of groups, in particular for amalgamated free products, has remained a challenging open problem.
@@ 1.2. Main results
In this paper we are able to make significant new progress and expand on the collection MF, PMF and notably PFF groups by developing a new $C^{*}$-algebraic approach to the problem. Our main result is the following:
@@
Suppose $G$ is an MF (resp., PMF, PFF) group and $H<G$ is separable (i.e., $H=\cap_{i\in\mathbb{N}}H_{i}$ where $H_{i}<G$ is a decreasing sequence of finite index subgroups). Let $L$ be a residually finite MF (resp., PMF, PFF) group such that either $G$ or $L$ is exact. Then the amalgamated free product $G*_{H}(H\times L)$ is MF (resp., PMF, PFF).
Our approach is inspired by recent progress on selflessness which is a form of intrinsic $C^{*}$-free independence in ultrapowers [42, 2, 38], in particular the work of Ozawa (see Section 4.6 of [10]). In order to demonstrate the power of our result, we quickly assemble a few corollaries. Firstly, we observe that there is an isomorphism $G\ast_{H}(H\times\mathbb{Z})\cong\ast_{H}G\rtimes\mathbb{Z}$ where $_{H}G$ denotes the infinite group double of $G$ over $H$, indexed by elements of $\mathbb{Z}$, and $\mathbb{Z}$ acts on it by permuting the copies of $G$ in accordance with the left multiplication action of $\mathbb{Z}$ on itself. We also recall that, if $G$ and $H$ are PFF and one of them is exact, then $G\times H$ is PFF (Lemma 2.4). PFF groups are also automatically residually finite. This yields:
@@
We point out that strikingly even the case of free products remained open for the PFF property until our work. Outside of families of LERF groups (see a list of such examples in [19]), a nice instance of such inclusions $H<G$ which can be fed into our Corollary 1.2 is quasi-convex subgroups of cubulated hyperbolic groups. Indeed, cubulated hyperbolic groups are subgroups of RAAGs [1] and hence PFF (our Corollary 1.3 below), and quasi-convex subgroups are separable [23]. Another natural setting for separability is in the world of graph products [30]. The following is our next main Corollary:
@@
Our result above offers a significant generalization of the breakthrough work of Magee–Thomas [36]. We emphasize that this combination result yields many new examples of MF and PMF groups, and even in the case of RAAGs, our proof of PMF is new and succinct. More importantly, the reach of our results is higher and can handle the PFF property for RAAGs and therefore all groups that embed virtually in them. For fundamental groups of closed hyperbolic 3-manifolds, this has been an open problem in the field [^1] We thank Ramon van Handel for sharing with us the statement of this open problem., which is now settled. This particular result of ours is of geometric significance because PFF (for free and surface groups) is exactly the ingredient that goes into the breakthrough work of A. Song on minimal surfaces in Euclidean unit spheres [48]. A. Song’s paper accesses PFF via the PPF result of Bordenave–Collins [7], and the follow up paper [3] accesses PFF via Louder–Magee’s PPF for surface groups [34]. However, our results open the door to large new families of groups without having to go through PPF, which has obstructions as we shall describe soon. This naturally suggests generalizations of the works [48, 3]. We document the following below (see [1, 23] and also [4, Theorem 5.27]):
@@
Let $N$ be a non positively curved, compact, orientable, aspherical 3-manifold with possibly empty boundary. Then $\pi_{1}(N)$ is PFF. More generally, all virtually special groups are PFF.
1.3. Comments to the reader
Optimality
@@ 1.3. Comments to the reader
We would like to first discuss optimality aspects of our work. First note that exactness is a rather essential ingredient, in for instance Corollary 1.3. Indeed even in the tensor product case it is not known if MF is preserved in full generality [21]. A natural question arises of whether our results can extend to accessing PPF. We point out that this is in fact not possible. For instance, our Corollary 1.3 simply cannot generalize to PPF as stated because as mentioned earlier, Magee has shown that the group $\mathbb{F}_{2}\times\mathbb{F}_{2}\times\mathbb{F}_{2}$ is not PPF [37], and this is indeed a RAAG. For this reason PFF seems to be the best one can hope for in these general families of groups. The challenge with PPF is that it is not amenable to being stable under direct products (or finite index overgroups), due to the following fundamental issue: tensoring does not preserve the image being in the orthogonal complement of the invariant vectors for permutation matrices. Another point we would like to make is that it is at the moment an intractable open problem whether soficity or Connes embeddability is closed under amalgamated free products. In fact, considering the examples in Theorem 1.1, it is unclear whether the separability assumption can be dropped, simply because even Connes embeddability is not known in that generality.
Concreteness of matrix models
At first glance, it may appear that our work reveals strongly convergent matrix models in a more abstract or existential sense. This is far from true. In fact, as we point out in the statement of Lemma 2.3, we arrive at a very concrete and natural sequence of group homomorphisms
which sends $g\in G$ to $(gK_{i},g)$ and $n\in\mathbb{Z}$ to $(n,e)$, and argue that in fact these maps in the ultraproduct will extend to the reduced $C^{*}$-algebra. Note that the groups on the right hand side are products of virtually free groups (for a very concrete reason, see Lemma 2.5) with $G$. Virtually free groups have a concrete strongly converging PFF model via [7] and an induced representation argument (see Lemma 2.4), and additionally $G$ is assumed to have a concrete PFF model. Combining these, we do in fact get concrete strongly converging models in our results, and in specific examples they can be tracked down explicitly.
A comment on bootstrapping
@@
We include a minor remark that our work also provides an alternative proof of PFF for limit groups also (see Remark 2.6). It is important to clarify however that the work [34] proves PPF for limit groups and this cannot be addressed currently by our method. The proof of this in our work combines our Theorem 1.1 and the fact that limit groups arise as iterated extensions of centralizers (Theorem 4.2 of [11]), and Hall’s theorem for limit groups [57]. We also point out for the benefit of the reader that the proof of Corollary 1.3 follows from the fact that every graph product is constructed in an iterated fashion using the amalgamated free products in Theorem 1.1, and additionally the amalgams are separable because they are retracts (Lemma 3.9 of [30]). We also remark that our approach might also naturally be applied to get more examples of MF/PMF/PFF groups among the family of graph wreath products [17].
Insights into the proof
A key accomplishment of our work is to isolate a precise and potent connection between some of the deep tools going into the emerging program on selflessness in $C^{*}$-algebras, and strong convergence of unitary representations. In effect, our approach is able to entirely reduce the problem of constructing strongly converging matrix models of groups of the form $G*_{H}(H\times\mathbb{Z})$, to just the case of free groups $\mathbb{F}_{n}$ ([20, 7]) and elementary constructions such as induced representations and tensor models via exactness. On one hand, handling such amalgamated free product groups allows us to cover large families at the same time thanks to beautiful results in geometric group theory around separability, RAAGs and virtual specialness [1, 23, 57, 30]. On the other hand, this construction is precisely what allows us access to the deep $C^{*}$-correspondence theory machinery (see Theorem 2.2), in order to upgrade our maps from just the groups to the reduced $C^{*}$-algebras.
Firstly, our maps from Lemma 2.3 are built into an ultraproduct of amalgamated free products (by passing to normal cores from the separability assumptions) which are each by design virtually free, and therefore admit various strongly converging models. In order for our maps to preserve norms, we use a Fell’s absorption trick by tensoring with a copy of $G$ throughout the procedure. Interestingly, this is the reason why we need exactness in the case that we replace $\mathbb{Z}$ with an arbitrary residually finite group $L$, and also the reason why one cannot automatically use these arguments to build PPF models. Continuing with the argument, we crucially use a remarkable isomorphism theorem (Theorem 2.2), which says that a universal Toeplitz–Pimnser type algebra satisfying just some simple orthogonality conditions, is automatically isomorphic to a reduced amalgamated free product of the type we have been considering. All we need to do is check the prescribed equalities with the “creation operators”, and we are then in business for strong convergence. We strongly urge the reader who is not familiar with these $C^{*}$-algebraic considerations to consider reading our proof in parallel with Section 4.6 of [10], which has a very neat treatment of the breakthroughs in $C^{*}$-correspondence theory around Fowler–Muhly–Raeburn’s gauge invariant-uniqueness theorem [16] and its consequences including the isomorphism theorem we use here.
@@ 1.4. Acknowledgments
1.5. Open Access and Data Statement
@@ 2. Proof of main result
Let $B\subset A$ be an inclusion of $C^{*}$-algebras, and $E:A\to B$ be a faithful conditional expectation. Then define the Toeplitz–Pimsner algebra of $E$ to be the universal $C^{*}$-algebra
@@
$T$ is called the creation operator. We recall that there is a canonical non-degenerate conditional expectation $E^{\prime}:\mathcal{T}(E)\to A$ satisfying $E^{\prime}(a_{1}Tb_{1}b_{2}^{\ast}T^{\ast}a_{2}^{\ast})=0$ for all $a_{i},b_{i}\in A$ [10, Theorem 4.6.6(2)].
A special case of the above is the classical Toeplitz algebra, which we will denote by $\mathcal{T}$. It corresponds to the case where $A=B=\mathbb{C}$ and $E=\text{id}_{\mathbb{C}}$. An alternative description is that $\mathcal{T}$ is the universal $C^{\ast}$-algebra of an isometry. In this case, the canonical conditional expectation yields a non-degenerated state $\omega$ on $\mathcal{T}$, which is called the vacuum state, and we shall always understand $\mathcal{T}$ as equipped with this state.
The following Theorem 2.2 is due to Shlyakhtenko in [46] in the von Neumann algebra setting (see also [16]). The proof in the $C^{*}$-setting is more or less identical. Nevertheless, we provide a proof for the convenience of the reader (following [10] wherein this result also appears). Our inspiration for employing this result in our context comes from the recent outburst of developments surrounding selfless $C^{*}$-algebras ([42, 2, 41, 27, 38, 24, 32, 33, 51, 52, 15, 5, 8, 29, 58, 14, 18]), especially the proof technique of Ozawa in [38] and also the upcoming key work [32]. The second and fourth authors are grateful to M. Junge and L. Robert for the collaboration on [32].
Before stating and proving Theorem 2.2, we emphasize that the beautiful gauge-invariant uniqueness theorem (this is explicitly [10, Theorem 4.6.18]) is used in the proof of Theorem 2.2. We highly encourage the reader to look at Section 4.6 in the treatise [10] which offers a clean treatment of these aspects.
@@
First consider $A\ast_{B}(B\otimes\mathcal{T})$ where $\mathcal{T}$ is the classical Toeplitz algebra. Let $S$ be the generating isometry in $\mathcal{T}$. Then for $a\in A$, $S^{*}aS=E(a)$. Hence there is a map $\pi:\mathcal{T}(E)\to A\ast_{B}(B\otimes\mathcal{T})$ which restricts to identity on $A$ and sends $T\mapsto S$. This map is surjective as $A\ast_{B}(B\otimes\mathcal{T})$ is generated by $A$ and $\mathcal{T}$.
We will prove $\pi$ is injective using the gauge-invariant uniqueness theorem for Toeplitz–Pimsner algebras [10, Theorem 4.6.18] (this is originally due to [16]). For this, note that the canonical conditional expectation $E^{\prime}:A\ast_{B}(B\otimes\mathcal{T})\to A$ acts by $\operatorname{id}\otimes\omega$ on the second factor, where $\omega$ is the vacuum state on the classical Toeplitz algebra. So we have, for $a\in A$,
@@
where we used that $S$ commutes with $E(a)\in B$. Thus, $\pi(A)\cap\overline{\operatorname{span}}\{a_{1}Sb_{1}b_{2}^{*}S^{*}a_{2}^{*}:a_{1},b_{1},b_{2},a_{2}\in A\}=\{0\}$, and hence $\pi$ satisfies the hypothesis of the gauge-invariant uniqueness theorem.
@@
which sends $g\in G$ to $(gK_{i},g)$ and $n\in\mathbb{Z}$ to $(n,e)$.
@@
Consider the sequence of group homomorphisms